OpenStudy (anonymous):

Find the probability that at least 4 of the 10 randomly selected parts weigh greater than 41 pounds. Mean 40 pounds SD 1.5 pounds

OpenStudy (perl):

Is that all the information given?

OpenStudy (anonymous):

yes

OpenStudy (perl):

Let's assume we are given the probability distribution of selecting a part. The mean of this distribution is 40 , and standard deviation is 1.5 pounds.

OpenStudy (perl):

It looks like there is more to this question. it says 'the parts' what parts?

OpenStudy (perl):

"Find the probability that at least 4 of the 10 randomly selected parts weigh greater than 41 pounds. " Was there a first part to this question

OpenStudy (anonymous):

Parts refers to what they are choosing not anything else. its like saying randomly selected fruits.

OpenStudy (perl):

The problem is, we need to know the probability of choosing one part that is greater than 41 pounds in order to answer the question. Can we assume the parts are normally distributed with a mean of 40 pounds and st. dev of 1.5?

OpenStudy (perl):

the original population of parts

OpenStudy (anonymous):

yes we can assume that

OpenStudy (anonymous):

i found the number for that was around 0.2525

OpenStudy (perl):

hmm, does it state that in the question (that's why i wasnt sure) also is this a multiple choice question, what are the choices .

OpenStudy (anonymous):

no it does not and its not multiple choice

OpenStudy (perl):

P ( X > 41) = normalcdf ( 41, 1e99, 40 , 1.5) = .25249 ~ .2525

OpenStudy (perl):

yeah i dont like to make assumptions that are not given explicitly :) this problem is a bit vague, but we can just impose our assumptions i guess

OpenStudy (perl):

ok so we know the probability of choosing 1 part that is greater than 41 pounds. now lets answer the question, suppose we sample 10 parts from the population. what is probability at least 4 is greater than 41 pounds

OpenStudy (anonymous):

yeah at this point I'm very ok with assumptions this has been giving me lots of trouble

OpenStudy (perl):

so we are looking at samples of ten (instead of samples of one)

OpenStudy (perl):

lets define a variable Y Y = number of parts which are success , or greater than 41 pounds note that Y can equal to 0,1,2,3,... 9,10

OpenStudy (anonymous):

ok

OpenStudy (perl):

the question is asking find P( Y>= 4) this is a binomial distribution problem

OpenStudy (perl):

it is easier to find the complement P( Y>= 4) = 1 - P( Y < 4 )

OpenStudy (anonymous):

ok

OpenStudy (perl):

1 - P( Y<4 ) = 1 - [ P( Y=0) + P( Y= 1) + P( Y = 2) + P( Y = 3 ) ]

OpenStudy (perl):

we need to find P( Y=0) + P( Y= 1) + P( Y = 2) + P( Y = 3 ) lets do this in separate calculations P(Y=0) = (1-.2525)^10 P(Y=1) = (10 choose 1 ) * (.2525)^1 * (1-.2525)^9 P(Y=2) = (10 choose 2 ) * (.2525)^2 *(1-.2525)^8 P(Y=3) = (10 choose 3 ) * (.2525)^3 *(1-.2525)^7

OpenStudy (perl):

if you want to be consistent you can write P (Y = 0) = (10 choose 0 ) * (.2525)^0 * ( 1-.2525)^10 , = 1 * 1 * (1-.2525)^10

OpenStudy (anonymous):

ok

OpenStudy (perl):

also you can do this on a TI 83 calculator , 1- P ( Y < 4) = 1 - P( Y <= 3 ) = 1 - binomcdf( 10, .2525, 3 ) = .22999

OpenStudy (anonymous):

can you elaborate on the calculator way??

OpenStudy (perl):

There is a built-in command binomcdf (binomial cumulative density function) that can be used to quickly determine "at most". Because this is a "cumulative" function, it will find the sum of all of the probabilities up to, and including. the command is 2ND -> VARS ->(press down until you see binomcdf ( binomcdf (number of trials, probability of occurrence, number of specific events) or shortened we have binomcdf (n, p, r)

OpenStudy (perl):

P(Y <= 3) , p = .2525, n = 10 is equal to binomcdf ( 10, .2525, 3 )

OpenStudy (perl):

also I used the fact that P ( Y<4 ) = P ( Y <= 3 ) , which makes sense since 0,1,2,3 are less than 4

OpenStudy (anonymous):

Woah i had no idea this was possible on the calc

OpenStudy (perl):

yes its nice

OpenStudy (perl):

there is a limitation though, for example if you want P ( 2<= Y <= 5 ) , you have to do P ( Y < = 5) - P ( Y <= 1 ) since P( Y<= r ) = binomcdf ( n, p , r )

OpenStudy (anonymous):

ah ok

OpenStudy (anonymous):

so whats next to solve this completely

OpenStudy (perl):

so you subtract that from 1 to get your answer

OpenStudy (perl):

1 - binomcdf( 10, .2525, 3 ) = .22999

OpenStudy (anonymous):

oh ok that makes sense

OpenStudy (anonymous):

wait does this answer the original question??

OpenStudy (perl):

yes

OpenStudy (perl):

there were two parts to answering the question. first we had to find the probability of picking 1 item that had a weight greater than 41 pounds (ie picking samples of one). Then the second part, suppose instead of sampling one item you sampled ten at a time, what is probability that at least 4 of the items have weight greater than 41 pounds. we needed the first part in order to find p = .2525 for the binomial distribution

OpenStudy (anonymous):

do u still need help @rowerguy508 ?

OpenStudy (perl):

do you agree with this? so far

OpenStudy (anonymous):

who?

OpenStudy (perl):

you Noland

OpenStudy (anonymous):

oh yes i do

OpenStudy (perl):

I mean if you have a different solution,

OpenStudy (anonymous):

great work guys

OpenStudy (perl):

i wasnt sure because the directions were a bit vague